# From Description to Score: Can LLMs Quantify Vulnerabilities?

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 41st ACM/SIGAPP Symposium on Applied Computing; March 23–27, 2026; Thessaloniki, Greeceisbn: 979-X-XXXX-XXXX-X/26/03ccs: Security and privacy Software and application securityccs: Security and privacy Vulnerability managementccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
Vulnerability management is a fundamental component of software security programs across organizations. Public databases such as the National Vulnerability Database (NVD), maintained by NIST’s Information Technology Laboratory (ITL) (NVD, 18), catalog and score newly disclosed vulnerabilities. In 2024, the NVD published 40,009 CVEs, over 38% more than in 2023 (gamblin2025cvedata, 9), increasing the demand for timely and consistent CVSS scoring (FIRST-CVSS, 8), which organizations rely on to prioritize remediation and allocate security resources.
This rapid growth has placed substantial strain on maintainers and threat intelligence providers. Throughout 2024, the NVD faced significant processing backlogs (munshaw2024nvd, 16), leaving many vulnerabilities unscored or inconsistently characterized and limiting organizations’ ability to assess risk effectively. Internally discovered vulnerabilities pose similar challenges, as they often lack standardized severity ratings, making prioritization subjective and resource-intensive.
On the other hand, recent advances in GenAI have shown that LLMs possess capabilities that go well beyond natural language generation. In particular, they have proven effective in classification tasks, including multi-class scenarios, by reframing classification as a text-to-text task (gretz2023zero, 10, 34). This allows LLMs to assign class labels based on textual input without requiring modifications to their underlying architecture (brown2020language, 2, 25). These advances prompt a natural question: to what extent can generative AI be leveraged to classify vulnerability metrics, such as those defined in the CVE system, using only the textual description of the vulnerability. If effective, such an approach could enable automated vulnerability scoring and help mitigate issues like the backlog observed in the NVD. Motivated by this question, the main contributions of this paper are as follows:
@@ 2.1. Common Vulnerability Scoring System
Figure 1. CVSS Metric Groups (cvssv3.1, 7)
@@ 2.2. Challenges of Triaging CVEs by NVD
The NVD has faced challenges in handling the surge of reported vulnerabilities, 28,818 CVEs in 2023 (gamblin2025cvedata, 9), overwhelming its 21-person team (cybersecuritydive2024cveanalysis, 11). By mid-2024, fewer than 10% of 12,700 CVEs since February had been analyzed (cipollone2024nvdcrisis, 3), delaying severity scores and product mappings vital to risk assessments (vaughannichols2024nvdslowdown, 30). Contributing factors include:
@@
  • Schema migration: Transition to JSON schema stalled pipelines (nist2024nvdnews, 20).
@@
These challenges expose systemic weaknesses. While funding and coordination are needed, modernizing and automating NVD processes, including automated vulnerability scoring, is critical to restoring timely vulnerability triage (therecord2024nistbacklog, 27).
@@ 3. Related Work
Challenges of Vulnerability Scoring. Manual vulnerability scoring is prone to inconsistency, as shown by Wunder et al. (wunder2024shedding, 31), who found that 68% of participants changed their CVSS ratings when reassessing the same vulnerabilities. Key metrics like Attack Vector and Privileges Required were especially error-prone. These results highlight the subjectivity of human scoring and support the need for automated approaches, such as those based on GenAI, to ensure more consistent and scalable assessments. Stoker et al. (threatt2024some, 28) showed that despite long-term trends in CVE data, CVSS ratings remained clustered around medium severity, often failing to reflect contextual risk. Similarly, Stoker et al. (nichols2021heuristic, 19) found inconsistencies in how vendors communicate risk. These findings motivate our use of GenAI to deliver standardized and consistent scoring.
ML and GenAI-based Scoring of Vulnerabilities. Recent studies have explored ML for creating explainable AI to help Counting Number Authority scoring vulnerabilities faster and more accurately. Manai et al. (manai2024helping, 12) used XGBoost with SHAP to produce interpretable CVSS v3.1 predictions. Yazigi (yazigi2025modernizing, 33) proposed a GenAI system that aggregates threat data for dynamic risk scoring. Alam et al. (alam2024ctibench, 1) introduced CTIBench, which benchmarks LLMs on five cybersecurity tasks, including CTI-VSP, closely aligned with our focus. Using zero-shot prompting on 1,000 CVEs, they found performance was sensitive to description length. Our study extends this by analyzing  31,000+ CVEs, evaluating class imbalance, inter-metric correlations, and model performance using precision, F1-score, and mean absolute error (MAE). Ullah et al. (ullah2024secLLMHolmes, 29) conducted a comprehensive evaluation of LLMs for vulnerability detection using their framework SecLLMHolmes. They evaluated eight LLMs, showing that even advanced models exhibit high false positive rates, inconsistent outputs, and fragile reasoning that breaks under minor code augmentations. Their findings highlight that LLMs are not yet robust enough for automated vulnerability detection in real-world settings, underscoring the need for further research into model reliability and reasoning faithfulness. It should be emphasized that the incorporation of CVE identifiers in prompts to LLMs, together with the corresponding description as in prior work (marchiori2025can, 13), allows models to retrieve known vectors and undermines true classification.
@@ 4.2. Data Collection
The study utilizes the publicly available CVE List maintained by MITRE (cve2025list, 24) as its primary data source. While all CVEs from 2004 to 2024 were initially retrieved, the analysis focuses on entries published from 2019 onward to ensure consistency with the CVSS v3.1 standard. After filtering for CVEs with complete CVSS v3.1 base metrics and valid English descriptions, the final dataset consists of 31,000+ entries. Each CVE, stored in JSON format, includes a human-readable description and eight CVSS v3.1 base metrics. Additional quality filters excluded CVEs with missing, non-English, or incomplete data. For the purposes of this study, the v3.1 CVSS base metrics and descriptions obtained from the CVE List are treated as ground truth. Although there may be some variance in scores and descriptions, the MITRE-maintained CVE List is a widely adopted and standardized source of vulnerability information (cve_process, 17).
@@ 4.3. Large Language Models
This study employs six models: GPT-4o (G4) (openai2024gpt4o, 22), GPT-5 (G5) (openai2025gpt5, 23), Llama-3.3-70B-Instruct (L) (meta2024llama3, 14), Gemini-2.5-Flash (GM) (deepmind2025geminiflash, 4), DeepSeek-R1 (DS) (deepseek2025r1, 5), and Grok-3 (GR) (xai2025grok3, 32) to predict CVSS base metrics directly from CVE descriptions. We intentionally selected a mix of high-performing closed-source and open-weight models to capture complementary strengths: closed models such as GPTs typically offer state-of-the-art accuracy, robustness, and safety features, while open models such as Llama provide transparency, reproducibility, and adaptability for security research. These characteristics make them strong candidates for CVSS scoring, which requires both precision and interpretability. We used Azure’s AI Foundry for GPT-4o, GPT-5, Llama, DeepSeek, and Grok, and Google’s AI Studio for Gemini.
@@ 4.4. Prompt Engineering
We followed OpenAI’s best practices (openai2025prompt, 21) in designing the prompt that directed the LLMs to generate CVSS base metric scores. One best practice is to be specific on the task and avoid combining results (mirzadeh2024gsm, 15). Therefore, our prompt did not ask for the overall CVSS score and instead asked for itemized scoring of metrics. We instructed the model to act as a cybersecurity expert. The prompt followed a two-step structure: (1) extract the eight base metrics, and (2) output them in a fixed format for consistency. To ensure deterministic outputs, if configurable, the temperature was set to 0, following CTIBench guidance (alam2024ctibench, 1). We evaluated zero-, two-, five-, and ten-shot prompts on a subset of 6,000 CVE entries from 2024. The two-shot variant achieved the best performance; therefore, it was adopted for all subsequent experiments. A Python script automated API calls in batches of 20 CVEs, appending predicted metrics to CSV outputs. Using this pipeline, we processed over 31,000 CVE records for downstream analysis.
@@ 6.2. Analysis of Individual LLM Classifiers
Attack Vector. Gemini led with 89.42% accuracy, surpassing the 72.65% baseline and achieving balanced scores (P=0.82, R=0.83, F1=0.82). GPT-5 followed at 87.96% (F1=0.88), showing the strongest recall (0.88) but slightly lower precision. GPT-4o reached 86.57%, also well above baseline with consistent metrics (F1 = 0.86). Llama (79.31%) and DeepSeek (79.01%) showed moderate results (F1=0.76–0.79), while Grok (83.00%) matched Gemini’s stability (F1 = 0.82). Although all models outperformed the baseline, multiple showed a bias toward the majority class, often predicting NETWORK over minority classes like LOCAL.
@@ 6.3. Analysis of the content of “description” field
Information Contents. We computed the information content (IC) of each description using the Semantic Correlation formula as proposed in prior work on intrinsic IC measures (seco2004intrinsic, 26). However, we found no strong correlation between IC and the model’s precision, suggesting that IC alone is not a determining factor in predictive performance.
@@ 9. Conclusion
This study evaluated the feasibility of using general-purpose LLMs, including GPT-4o, GPT-5, Llama-3.3-70B-Instruct, Gemini-2.5-Flash, DeepSeek-R1, and Grok-3, to automate CVSS v3.1 base metric classification from vulnerability descriptions. Across more than 31,000 CVEs, GPT-5 and Gemini-2.5-Flash showed the strongest overall performance, particularly for metrics such as Attack Vector and User Interaction. However, all models struggled with minority classes and were sensitive to class imbalance and ambiguous descriptions. The meta-classifier offered only modest accuracy improvements, indicating that fine-tuning and domain-specific enhancements remain necessary. Overall, our findings show that LLMs can support scalable vulnerability triage but also reveal key limitations that must be addressed. Future work should explore instruction tuning and the integration of external contextual signals to further improve prediction reliability in operational settings. All project code and data have been released as open-source materials (ecdeans_AI-Vuln-Management, 6).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 41st ACM/SIGAPP Symposium on Applied Computing; March 23–27, 2026; Thessaloniki, Greeceisbn: 979-X-XXXX-XXXX-X/26/03journalyear: 2026copyright: ccconference: The 41st ACM/SIGAPP Symposium on Applied Computing; March 23–27, 2026; Thessaloniki, Greecebooktitle: The 41st ACM/SIGAPP Symposium on Applied Computing (SAC ’26), March 23–27, 2026, Thessaloniki, Greecedoi: 10.1145/3748522.3779726isbn: 979-8-4007-2294-3/2026/03ccs: Security and privacy Software and application securityccs: Security and privacy Vulnerability managementccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
Vulnerability management is a fundamental component of software security programs across organizations. Public databases such as the National Vulnerability Database (NVD), maintained by NIST’s Information Technology Laboratory (ITL) (undefq, ), catalog and score newly disclosed vulnerabilities. In 2024, the NVD published 40,009 CVEs, over 38% more than in 2023 (undefh, ), increasing the demand for timely and consistent CVSS scoring (undefg, ), which organizations rely on to prioritize remediation and allocate security resources.
This rapid growth has placed substantial strain on maintainers and threat intelligence providers. Throughout 2024, the NVD faced significant processing backlogs (undefo, ), leaving many vulnerabilities unscored or inconsistently characterized and limiting organizations’ ability to assess risk effectively. Internally discovered vulnerabilities pose similar challenges, as they often lack standardized severity ratings, making prioritization subjective and resource-intensive.
On the other hand, recent advances in GenAI have shown that LLMs possess capabilities that go well beyond natural language generation. In particular, they have proven effective in classification tasks, including multi-class scenarios, by reframing classification as a text-to-text task (undefi, ; undefag, ). This allows LLMs to assign class labels based on textual input without requiring modifications to their underlying architecture (undefa, ; undefx, ). These advances prompt a natural question: to what extent can generative AI be leveraged to classify vulnerability metrics, such as those defined in the CVE system, using only the textual description of the vulnerability. If effective, such an approach could enable automated vulnerability scoring and help mitigate issues like the backlog observed in the NVD. Motivated by this question, the main contributions of this paper are as follows:
@@ 2.1. Common Vulnerability Scoring System
Figure 1. CVSS Metric Groups (undeff, )
@@ 2.2. Challenges of Triaging CVEs by NVD
The NVD has faced challenges in handling the surge of reported vulnerabilities, 28,818 CVEs in 2023 (undefh, ), overwhelming its 21-person team (undefj, ). By mid-2024, fewer than 10% of 12,700 CVEs since February had been analyzed (undefb, ), delaying severity scores and product mappings vital to risk assessments (undefac, ). Contributing factors include:
@@
  • Schema migration: Transition to JSON schema stalled pipelines (undefs, ).
@@
These challenges expose systemic weaknesses. While funding and coordination are needed, modernizing and automating NVD processes, including automated vulnerability scoring, is critical to restoring timely vulnerability triage (undefz, ).
@@ 3. Related Work
Challenges of Vulnerability Scoring. Manual vulnerability scoring is prone to inconsistency, as shown by Wunder et al. (undefad, ), who found that 68% of participants changed their CVSS ratings when reassessing the same vulnerabilities. Key metrics like Attack Vector and Privileges Required were especially error-prone. These results highlight the subjectivity of human scoring and support the need for automated approaches, such as those based on GenAI, to ensure more consistent and scalable assessments. Stoker et al. (undefaa, ) showed that despite long-term trends in CVE data, CVSS ratings remained clustered around medium severity, often failing to reflect contextual risk. Similarly, Stoker et al. (undefr, ) found inconsistencies in how vendors communicate risk. These findings motivate our use of GenAI to deliver standardized and consistent scoring.
ML and GenAI-based Scoring of Vulnerabilities. Recent studies have explored ML for creating explainable AI to help Counting Number Authority scoring vulnerabilities faster and more accurately. Manai et al. (undefk, ) used XGBoost with SHAP to produce interpretable CVSS v3.1 predictions. Yazigi (undefaf, ) proposed a GenAI system that aggregates threat data for dynamic risk scoring. Alam et al. (undef, ) introduced CTIBench, which benchmarks LLMs on five cybersecurity tasks, including CTI-VSP, closely aligned with our focus. Using zero-shot prompting on 1,000 CVEs, they found performance was sensitive to description length. Our study extends this by analyzing  31,000+ CVEs, evaluating class imbalance, inter-metric correlations, and model performance using precision, F1-score, and mean absolute error (MAE). Ullah et al. (undefab, ) conducted a comprehensive evaluation of LLMs for vulnerability detection using their framework SecLLMHolmes. They evaluated eight LLMs, showing that even advanced models exhibit high false positive rates, inconsistent outputs, and fragile reasoning that breaks under minor code augmentations. Their findings highlight that LLMs are not yet robust enough for automated vulnerability detection in real-world settings, underscoring the need for further research into model reliability and reasoning faithfulness. It should be emphasized that the incorporation of CVE identifiers in prompts to LLMs, together with the corresponding description as in prior work (undefl, ), allows models to retrieve known vectors and undermines true classification.
@@ 4.2. Data Collection
The study utilizes the publicly available CVE List maintained by MITRE (undefw, ) as its primary data source. While all CVEs from 2004 to 2024 were initially retrieved, the analysis focuses on entries published from 2019 onward to ensure consistency with the CVSS v3.1 standard. After filtering for CVEs with complete CVSS v3.1 base metrics and valid English descriptions, the final dataset consists of 31,000+ entries. Each CVE, stored in JSON format, includes a human-readable description and eight CVSS v3.1 base metrics. Additional quality filters excluded CVEs with missing, non-English, or incomplete data. For the purposes of this study, the v3.1 CVSS base metrics and descriptions obtained from the CVE List are treated as ground truth. Although there may be some variance in scores and descriptions, the MITRE-maintained CVE List is a widely adopted and standardized source of vulnerability information (undefp, ).
@@ 4.3. Large Language Models
This study employs six models: GPT-4o (G4) (undefu, ), GPT-5 (G5) (undefv, ), Llama-3.3-70B-Instruct (L) (undefm, ), Gemini-2.5-Flash (GM) (undefd, ), DeepSeek-R1 (DS) (undefe, ), and Grok-3 (GR) (undefae, ) to predict CVSS base metrics directly from CVE descriptions. We intentionally selected a mix of high-performing closed-source and open-weight models to capture complementary strengths: closed models such as GPTs typically offer state-of-the-art accuracy, robustness, and safety features, while open models such as Llama provide transparency, reproducibility, and adaptability for security research. These characteristics make them strong candidates for CVSS scoring, which requires both precision and interpretability. We used Azure’s AI Foundry for GPT-4o, GPT-5, Llama, DeepSeek, and Grok, and Google’s AI Studio for Gemini.
@@ 4.4. Prompt Engineering
We followed OpenAI’s best practices (undeft, ) in designing the prompt that directed the LLMs to generate CVSS base metric scores. One best practice is to be specific on the task and avoid combining results (undefn, ). Therefore, our prompt did not ask for the overall CVSS score and instead asked for itemized scoring of metrics. We instructed the model to act as a cybersecurity expert. The prompt followed a two-step structure: (1) extract the eight base metrics, and (2) output them in a fixed format for consistency. To ensure deterministic outputs, if configurable, the temperature was set to 0, following CTIBench guidance (undef, ). We evaluated zero-, two-, five-, and ten-shot prompts on a subset of 6,000 CVE entries from 2024. The two-shot variant achieved the best performance; therefore, it was adopted for all subsequent experiments. A Python script automated API calls in batches of 20 CVEs, appending predicted metrics to CSV outputs. Using this pipeline, we processed over 31,000 CVE records for downstream analysis.
@@ 6.2. Analysis of Individual LLM Classifiers
Attack Vector. Gemini led with 89.42% accuracy, surpassing the 72.65% baseline and achieving balanced scores (P=0.82, R=0.83, F1=0.82). GPT-5 followed at 87.96% (F1=0.88), showing the strongest recall (0.88) but slightly lower precision. GPT-4o reached 86.57%, also well above baseline with consistent metrics (F1 = 0.86). Llama (79.31%) and DeepSeek (79.01%) showed moderate results (F1=0.760.79)(F1=0.76–0.79), while Grok (83.00%) matched Gemini’s stability (F1 = 0.82). Although all models outperformed the baseline, multiple showed a bias toward the majority class, often predicting NETWORK over minority classes like LOCAL.
@@ 6.3. Analysis of the content of “description” field
Information Contents. We computed the information content (IC) of each description using the Semantic Correlation formula as proposed in prior work on intrinsic IC measures (undefy, ). However, we found no strong correlation between IC and the model’s precision, suggesting that IC alone is not a determining factor in predictive performance.
@@ 9. Conclusion
This study evaluated the feasibility of using general-purpose LLMs, including GPT-4o, GPT-5, Llama-3.3-70B-Instruct, Gemini-2.5-Flash, DeepSeek-R1, and Grok-3, to automate CVSS v3.1 base metric classification from vulnerability descriptions. Across more than 31,000 CVEs, GPT-5 and Gemini-2.5-Flash showed the strongest overall performance, particularly for metrics such as Attack Vector and User Interaction. However, all models struggled with minority classes and were sensitive to class imbalance and ambiguous descriptions. The meta-classifier offered only modest accuracy improvements, indicating that fine-tuning and domain-specific enhancements remain necessary. Overall, our findings show that LLMs can support scalable vulnerability triage but also reveal key limitations that must be addressed. Future work should explore instruction tuning and the integration of external contextual signals to further improve prediction reliability in operational settings. All project code and data have been released as open-source materials (undefc, ).